OpenStudy (anonymous):

if p+pq is 4 times p-pq, which of the following has exactly one value? (pq≠0) a) p b) q c) pq d) p+pq e) p-pq

OpenStudy (anonymous):

\[p+pq=4(p-pq)\\ p+pq=4p-4pq\\ 5pq=3p\] Since \(pq\not=0\), you know that \(p\not=0\) and \(q\not=0\) (zero product property). So, you can divide both sides by \(p\): \[5q=3\]

OpenStudy (anonymous):

@SithsAndGiggles wait so what the answer because i dont know how you continue form 5q=3 unless its q for q=3/5

OpenStudy (anonymous):

So you get \(q=\dfrac{3}{5}\). Plug that into the original equation and solve for \(p\): \[p+\frac{3}{5}p=4\left(p-\frac{3}{5}p\right)\\ \frac{8}{5}p=\frac{8}{5}p\] You get an identity; this is true for all \(p\). It doesn't matter what you plug in for \(p\), as long as you have \(q=\dfrac{3}{5}\). This means \(p\) does not have exactly one value, and neither does any expression containing \(p\).

OpenStudy (anonymous):

so p means more than 1 value, so the equation without p is the answer?

OpenStudy (anonymous):

Yes, only \(q\) has exactly one value. Any "expression" (not equation) containing \(p\) would not have one value, since the equation above holds for all \(p\) and only one value of \(q\).

OpenStudy (anonymous):

ok thanx^^